# Simulating Prosocial Behavior and Social Contagion in LLM Agents under Institutional Interventions

### v1

@@ 5.1 Impact of Structural Inequity on Individual Prosocial Behavior
To assess whether structural inequity suppresses prosocial motivation, we simulate two common forms of policy asymmetry: reward asymmetry, where some agents receive recognition or benefits for prosocial actions while others do not, and burden asymmetry, where agents perform the same task but incur unequal burden based on income level. We measure changes in prosocial intention before and after asymmetry exposure on Deepseek-v3 and GPT-4o. Figure 8 shows that inequity leads to substantial reductions in prosociality. Under reward asymmetry, prosocial scores decline by 25–32%, while under burden asymmetry the reduction ranges from 19–23%. In both cases, the drop is more pronounced when inequity involves recognition rather than burden, suggesting that fairness in acknowledgment carries greater psychological weight for LLM agents. These findings indicate that LLMs are sensitive to structural unfairness, and that such asymmetries significantly reduce their willingness to engage in prosocial behavior.
5.2 Contagion Effects of Inequity in Social Networks
@@
We next examine whether the effects of structural inequity can propagate through social interactions. Using the GPT-4o model, we simulate the diffusion of inequity in a small-world network of 104 agents over 30 iterations. Initially, 20% of agents are randomly assigned to experience burden asymmetry. In each round, 10% of the network edges are activated, allowing agents to observe the behavior of their active neighbors and update their own prosocial tendencies accordingly.
@@
Figure 8 visualizes the contagion effects of policy inequity. The top panel tracks the spread of perceived unfairness over time. An agent is marked as indirectly exposed to inequity if any of its activated neighbors has previously experienced unfair treatment. By iteration 10, the majority of agents, regardless of whether they were initially affected, begin to report elevated perceptions of unfairness. This suggests that structural inequity spreads through the network via social observation and inference. The bottom panel shows the corresponding change in prosocial behavior. Nodes with high prosociality, indicated by dark blue, steadily decline in number, while low-prosociality nodes, shown in light blue, become more prevalent. This trend reveals that reduced willingness to act prosocially is not confined to those directly impacted, but spreads throughout the network through behavioral contagion. These findings underscore the systemic consequences of policy-induced inequity. Localized unfair treatment can propagate through social connections and lead to widespread erosion of prosocial norms at the population level.
5.3 Attribution of Behavioral Decline to Perceived Unfairness
@@
To test whether the decline in prosocial behavior under inequitable policies is driven by agents’ internal fairness assessments, we ask each agent to rate their perceived unfairness at every iteration of the simulation. This score reflects the extent to which the agent feels unequally treated based on the observed conditions of its neighbors within the network. Figure 9 illustrates the joint temporal dynamics of perceived unfairness and prosocial tendency over 30 iterations. In both the reward and burden asymmetry conditions, we observe a consistent pattern: as perceived unfairness rises in early rounds and stabilizes at a high level, the average prosocial tendency concurrently declines and remains suppressed. Notably, this decline occurs even among agents not directly affected by policy inequity, indicating that fairness perception can propagate socially and shape collective norms. Insets show a strong negative correlation between perceived unfairness and prosocial scores across agents, supporting the psychological plausibility of fairness-based moral disengagement. In both conditions, agents who felt more unfairly treated were systematically less likely to engage in prosocial behaviors. These findings highlight perceived unfairness as a key explanatory mechanism behind the erosion of prosocial behavior, offering a cognitive account for how structural inequities undermine social cohesion.
6 Conclusion

### v2

@@ Introduction
To assess whether inequity suppresses prosocial motivation, we simulate two common forms of policy asymmetry: reward asymmetry, where some agents receive recognition or benefits for prosocial actions while others do not, and burden asymmetry, where agents perform the same task but incur unequal burden based on income level. We measure changes in prosocial intention before and after asymmetry exposure on Deepseek-v3 and GPT-4o. Figure 7 shows that inequity leads to substantial reductions in prosociality. Under reward asymmetry, prosocial scores decline by 25–32%, while under burden asymmetry the reduction ranges from 19–23%. In both cases, the drop is more pronounced when inequity involves recognition rather than burden, suggesting that fairness in acknowledgment carries greater psychological weight for LLM agents. These findings indicate that LLMs are sensitive to structural unfairness, and such asymmetries significantly reduce their probability to engage in prosocial behavior.
Contagion Effects of Inequity in Social Networks
@@
We next examine whether the effects of structural inequity can propagate through social interactions. Using the GPT-4o model, we simulate the diffusion of inequity in a small-world network of 104 agents over 30 iterations. Initially, 20% of agents are randomly assigned to experience burden asymmetry. In each round, 10% of the network edges are activated, allowing agents to observe the behavior of their active neighbors and update their prosocial tendencies.
@@
Figure 7 visualizes the contagion effects of policy inequity. The top panel tracks the spread of perceived unfairness over time. An agent is marked as indirectly exposed to inequity if any of its activated neighbors has previously experienced unfair treatment. By iteration 10, the majority of agents, regardless of whether they were initially affected, begin to report elevated perceptions of unfairness. This suggests that structural inequity spreads through the network via social observation and inference. The bottom panel shows the corresponding change in prosocial behavior. Nodes with high prosociality, indicated by dark blue, steadily decline in number, while low-prosociality nodes, shown in light blue, become more prevalent. This trend reveals that reduced prosociality is not confined to those directly impacted, but spreads throughout the network through behavioral contagion. These findings underscore the systemic consequences of policy-induced inequity. Localized unfair treatment can propagate through social connections and lead to widespread erosion of prosocial norms at the population level.
Attribution of Behavioral Decline to Unfairness
@@
To test whether the decline in prosocial behavior under inequitable policies is driven by agents’ internal fairness assessments, we ask each agent to rate their perceived unfairness at every iteration of the simulation. This score reflects the extent to which the agent feels unequally treated based on the observed conditions of its neighbors within the network. Figure 8 illustrates the joint temporal dynamics of perceived unfairness and prosocial tendency over 30 iterations. In both conditions, we observe a consistent pattern: as perceived unfairness rises in early rounds and stabilizes at a high level, the average prosocial tendency concurrently declines and remains suppressed. Notably, this decline occurs even among agents not directly affected by policy inequity, indicating that fairness perception can propagate socially and shape collective norms. Insets show a strong negative correlation between perceived unfairness and prosocial scores across agents, supporting the psychological plausibility of fairness-based moral disengagement. In both conditions, agents who felt more unfairly treated were systematically less likely to engage in prosocial behaviors. These findings highlight perceived unfairness as a key explanatory mechanism behind the erosion of prosocial behavior, offering a cognitive account for how structural inequities undermine social cohesion.
Conclusion